OpenStudy (anonymous):

The wattage rating of an appliance is given as W, in watts, and varies jointly as the resistance, R, in ohms, and as the square of the current, I, in amperes. If the wattage is 36 watts when the resistance is 100 ohms and the current is 0.6 amperes, find the wattage when the resistance is 300 ohms and the current is 0.5 ampere

OpenStudy (anonymous):

\[W = k \times (I^2R)\] You are given with W = 36, R = 100 and I = 0.6 Put the values and find k first..

OpenStudy (anonymous):

\[36 = k \times(100 \times (0.6)^2\] \[36k = 36\] Now tell me what is the value of k??

OpenStudy (anonymous):

@prabhranu 22 reply vi karna hunda hai es site te..

OpenStudy (anonymous):

\[\Huge \color{green} {Reply}\]

OpenStudy (anonymous):

thank